OpenStudy (anonymous):

Power series problem. (x-2)^n/(n^n) Find the radius and interval of convergence.

OpenStudy (aykayyy):

i actually just learned this today lol..... try the ratio test!

OpenStudy (anonymous):

I did, but am not sure if I am doing ti correctly

OpenStudy (anonymous):

I would do the nth root test?

OpenStudy (aykayyy):

\[\lim_{n \rightarrow \infty} \left| \frac{ a _{n+1} }{ a _{n} } \right|\]

OpenStudy (anonymous):